A cyclist moves with a constant speed of 25 m/s as it passes over the top of a hill whose radius of curvature is 126 m. The mass of the cyclist and her cycle has a mass of 350 kg. Determine the apperant weight of the cyclist at the top of the hill.
1
Expert's answer
2012-07-06T07:34:24-0400
Taking acceleration due to gravity g = 9.82 m/s² , radius r = 126 m, mass m = 350 kg we get centripetal force

F = mv²/r& = 350 * 625 / 126 = 1736.11 Newtons (N).

Apparent weight is

W =& m*g& - 1736.11 = 3437 - 1736.11 = 1700.89 N,

then dividing by 9.82 m/s² we get 173.207 kg.